# A candidate quadruple AGN system at

### v1

@@
Multiple galaxies hosting active galactic nuclei (AGNs) at kpc separation from each other are exceedingly rare, and in fact, only one quadruple AGN is known so far. These extreme density peaks are expected to pinpoint protocluster environments and therefore be surrounded by large galaxy overdensities. In this letter, we present another quadruple AGN candidate at including two SDSS quasars at a separation of roughly 480 kpc. The brighter quasar is accompanied by two AGN candidates (a type 1 AGN and a likely type 2 quasar) at close ( kpc) separation identified through emission line ratios, line widths and high ionization lines like N v . The extended Ly emission associated with the close triple system is more modest in extent and brightness compared to similar multiple AGN systems and could be caused by ram-pressure stripping of the type-2 quasar host during infall into the central dark matter halo. The projected evolution of the system into a galaxy cluster with the AGN host galaxies forming the brightest cluster galaxy needs to be further tested by galaxy overdensity studies on large scales around the quadruple AGN candidate. If confirmed as a quadruple AGN with X-ray observations or rest-frame optical line ratios, this system would represent the second AGN quartet, the highest-redshift multiplet and the closest high-redshift triplet known.
@@
Galaxies: clusters: general – Galaxies: evolution – quasars: general1 Introduction
@@
While single bright quasars may pinpoint protocluster cores, physically associated multiple AGN systems could be even more reliable tracers of local overdensities (Onoue et al., 2018; Sandrinelli et al., 2018), albeit also not a firm indicator (Fukugita et al., 2004; Green et al., 2011), as a number of highly active, massive galaxies require ample supply of cold gas feeding supermassive black holes (SMBHs) and star formation. In the framework of hierarchical structure formation, such objects are most likely to reside in peaks of the density field populating the rare massive end of the halo mass function (Bhowmick et al., 2020). In the absence of dense environments, multiple AGN systems would be exceedingly rare due to their transient nature (Hennawi et al., 2006).
Interactions of three SMBHs are also predicted to be almost always necessary to form ultra-massive black holes with black hole masses (Hoffman et al., 2023), further highlighting their importance in cluster formation as the central brightest cluster galaxies commonly host ultra-massive black holes (e.g., Dullo et al. 2017; Mehrgan et al. 2019).
So far, only a few multiple AGN systems on kiloparsec scales are known. These objects are likely in an early stage of galaxy merger and might share a dark matter halo, but not their host galaxies. In contrast to that, sub-arcsecond triple AGNs trace local late-stage mergers with multiple cores in the same host galaxy (e.g., Liu et al. 2011; Kalfountzou et al. 2017; Pfeifle et al. 2019). This work concentrates on the former AGN tuples on kiloparsec scales, of which only five are known.
Specifically, Hennawi et al. (2015) report the discovery of a quasar quartet at the peak of cosmic noon, . The AGN are embedded in a massive Ly nebula extending by more than 300 kpc, indicative of a substantial reservoir of cool ( K) gas (). The AGN are all found within a sphere of projected radius ¡150 kpc. The first triple quasar to be identified also resides at a redshift of 2 within a projected distance of less than 50 kpc, consistent with the typical onset of AGN activity in galaxy mergers (Djorgovski et al., 2007). Another triplet has been detected at , consisting of a close pair with a companion at 200 kpc projected distance (Farina et al., 2013). The third triplet known is likewise embedded in a massive Ly nebula at and comprises two quasars and one faint AGN at projected distances below 100 kpc (Arrigoni Battaia et al., 2018). A fourth triplet was found in a sample of almost 20,000 AGN candidates at a redshift of 1.13 (Assef et al., 2018).
Here, we report on the discovery of a quadruple AGN candidate at consisting of a close triplet, contained within 20 kpc in projection, and a forth quasar at a separation of 478 kpc.
Throughout this work, we assume a flat CDM cosmology with , and (Planck Collaboration et al., 2020). Reported quasar magnitudes are in the AB system.
@@ 2 Observation and Data Analysis
One quasar in the system, SDSS J101254.73+033548.7 (hereinafter QSO1), has been targeted with the Multi Unit Spectroscopic Explorer (MUSE; Bacon et al. 2010) on the Very Large Telescope as part of a survey to study the environment of quasar pairs (Herwig et al., 2024). QSO1 is a known member of a quasar pair from the SDSS DR12 (Pâris et al., 2017). The projected distance to the second quasar in the pair SDSS J101251.06+033616.5 (hereinafter QSO2) is 478 kpc. For QSO2, only the SDSS spectrum is available. Here we briefly describe the observations and data reduction. More detailed descriptions of sample selection, observation and subsequent data analysis can be found in Herwig et al. (2024). The MUSE data were acquired in Wide Field Mode on March 14, 2018 (seeing of ), yielding a field of view (FoV) of (pixel scale of ) and targeting only QSO1 as the separation of the pair is larger than the MUSE FoV. Three exposures were taken with on-source time each, rotated by 90 degrees with respect to each other and covering a wavelength range of with a sampling of . The data have been reduced following Farina et al. (2019) and González Lobos et al. (2023). Specifically, we use the MUSE pipeline v 2.8.7 (Weilbacher et al., 2014, 2016, 2020) to perform bias- and sky-subtraction, flat fields, twilight and illumination correction and wavelength and flux calibration. To remove residual sky emission lines, we use the Zurich Atmospheric Purge (Soto et al., 2016). After masking gaps between the CCDs and correcting offsets between the exposures with custom routines, the three exposures are median combined to obtain the final data cube. To take into account correlated noise introduced during data processing, the variance cube produced by the pipeline is rescaled layer-by-layer to the RMS of a background region in the science cube, resulting in a mean pixel-wise noise level at the position of QSO1 of .
@@ 3.1 Discovery of a candidate quadruple AGN
By collapsing the final MUSE cube along the wavelength axis, we obtain a continuum image of the field surrounding QSO1. In this image (r-band in Fig. 1, left), two additional continuum sources are visible in close proximity to QSO1. While similar geometries can be seen as a result of gravitational lensing of a single source, we can confidentily reject this possibility in this case (Appendix A). We instead identify both of these sources (hereinafter called AGNc1 and AGNc2) as AGN candidates as detailed in the following. To constrain their nature, we extract spectra from the MUSE cube within an aperture of 0.8 to minimize contamination from each other’s PSFs. To determine their positions, we first model the PSF of a star within the MUSE FoV with a 2D Moffat function and fit the model to the three continuum sources observed with MUSE (QSO1, AGNc1, AGNc2) by fixing the FWHM and leaving amplitude and position free to vary (Fig.1, middle panel). The peak of the Moffat function is used as center of the aperture for spectral extraction and is indicated with pink circles (crosses) in the middle (right) panel of Fig. 1. To reduce noise while retaining the signal, the spectra are smoothed by a 1D Gaussian kernel with pixels, corresponding to . Before smoothing the spectrum of AGNc2, a sky emission line at 6365.3 is masked as it would otherwise be convolved with the C iv line. The spectra are shown in Fig. 2, together with the SDSS spectrum of QSO2. The two QSOs and two AGN candidates are all detected in Ly, confirming that they reside at roughly the same redshift. QSO1 and AGNc1 both show a broad C iii] emission line, while in the spectrum of AGNc2, there is a tentative detection of a narrow C iii] line as well as a prominent narrow, blue-shifted C iv line. A He ii emission line is tentatively detected in all four objects. Within the aperture of used for spectral extraction, the -band magnitude for the three AGNs observed with MUSE is mag, mag and mag. In SDSS DR17 (Abdurro’uf et al., 2022), the -band magnitude of QSO2 is reported as . However, the SDSS fiber diameter of 3 is much larger than the aperture used in this work and consequently, the magnitude of QSO1 is reported to be much brighter in SDSS with . Between QSO1 and AGNc1, we measure an angular separation of 2.1, corresponding to the distance kpc at this redshift. QSO1 and AGNc2 are separated by 2.6 or 20 kpc, and QSO2 is 61.6 or 478 kpc away from QSO1. We obtain the systemic redshift by fitting the C iii] line with a double Gaussian model and correcting by the characteristic velocity shift of between the peak of the line complex and the systemic redshift for quasars (Shen et al., 2016). The value obtained for AGNc1 is likely overestimated as a redshift of implies a highly unusual Ly line shape that is completely dominated by the blue wing (Fig. 4). If this was the case, the He ii line might be misidentified and possibly not detected. Positions of the AGN candidates and projected distance between each other as well as redshift and velocity shifts are summarized in Table 1.
@@
We fit the spectra as described in Appendix B and summarize the resulting line fluxes, widths and upper flux limits in Table 2. Black hole mass estimates for QSO1 and QSO2 using the relation and absolute accuracy estimate derived for C iv in Vestergaard & Peterson (2006) are and . The line width of AGNc2 is narrow with and consistent with a Type II AGN, while the spectrum of AGNc1 is dominated by broad lines of indicative of a Type I AGN.
@@
At the position of QSO1, radio emission is detected with a flux density of in VLASS (Lacy et al., 2020) and in VLA FIRST (Becker et al., 1994), leading to a radio slope of (). Additionally using the UV continuum slope from the fit to the spectrum of QSO1 (Appendix B) of (), we can calculate the ratio (Kellermann et al., 1989) by extrapolating the rest-frame flux densities and obtain , classifying QSO1 as radio-loud. Although no additional radio emission is detected, we cannot exclude that AGNc1 and AGNc2 are radio-loud due to the shallow surveys depth and their dim UV continuum.
[…]
QSO1 | 10:12:54.74 | +03:35:48.93 | 3.166 0.001 | 0 | 0/ 0 | |
AGNc1 | 10:12:54.88 | +03:35:49.22 | 3.183 0.008 | -1590 766 | 2.1/16 | |
AGNc2 | 10:12:54.77 | +03:35:46.36 | 3.158 0.001 | 761 166 | 2.6/20 | |
[…]
3.1.1 Emission line diagnostics
A commonly adopted diagnostic diagram to distinguish AGNs from star forming galaxies (SFGs) in rest-frame UV uses the line ratios C iv /C iii] (C43), mainly influenced by the ionization parameter, and C iii]/He ii (C3He2), which is sensitive to metallicity and allows distinction between shocks and AGN photoionization (Feltre et al., 2016). Hirschmann et al. (2019) propose demarcation lines to distinguish between active and inactive galaxies, but current models are not capable of fully reproducing observed line ratios, leading to large impurities in the selection of AGNs and specifically faint AGNs. Classification is further complicated because certain types of galaxies would not appear in such a diagnostic diagram: high- radio galaxies (HzRGs) commonly have no detection of C iv (56 % in De Breuck et al. 2000) and SFGs usually show C iv in absorption instead of emission (e.g. Shapley et al. 2003; Calabrò et al. 2022). In Fig. 3 we show the line ratios of the candidate AGNs and SDSS quasars together with line ratios of SFGs and different types of AGNs from the literature. Both quasars and both AGN candidates as well as 30 % of the literature AGNs fall within the composite region of the diagram, indicating contribution from both star formation and AGN activity. Surprisingly, the line ratios of QSO1 indicate the highest contribution from star formation of all four objects although it is definitively classified as quasar, further highlighting the difficulty in AGN selection based on UV line ratios. It is also notable that the stacks of observed SFGs selected based on C iii] emission are classified as composite objects as well and show similar C3He2 ratios as AGNc1.
However, the detection of N v and the extreme line width of AGNc1 larger than firmly point towards an AGN origin. If these broad lines were produced by galactic-scale outflows instead, the violent kinematics should lead to a high contribution from shocks and likely lead to a detectable C iv emission line.
@@ 3.1.1 Emission line diagnostics
[…]
@@ 4.1 Clues from the peculiar extended Ly emission
The pink contours in Fig. 1 show the 2 and 5 levels of extended Ly emission, corresponding to a surface brightness of and respectively, and determined after subtracting empirical quasar point spread functions and continuum sources from the combined cube. This analysis is described in detail in Herwig et al. (2024) and Lobos et al. (2023). The nebular emission extends for about 50 kpc in the south-east direction with respect to QSO1, starting from the projected location of AGNc2. Additional Ly emission coincides with the position of AGNc1. Surprisingly, the surface brightness level of the extended Ly emission is much lower than typically found for multiple AGN systems (Hennawi et al., 2015; Arrigoni Battaia et al., 2018) and the AGN candidates are not embedded in the nebula. Instead, the flux-weighted centroid of the extended emission is offset from QSO1 by 34 kpc (Herwig et al., 2024). Possible explanations for this could be a faster evolution and therefore early cool gas depletion of outlier density peaks in the universe or a later merger stage of the close AGN triplet candidate presented in this work. Moreover, strong feedback effects might have already depleted the circumgalactic medium of this system of cool gas.

### v2

@@
Multiple galaxies hosting active galactic nuclei (AGNs) at kpc separation from each other are exceedingly rare, and in fact, only one quadruple AGN is known so far. These extreme density peaks are expected to pinpoint protocluster environments and therefore be surrounded by large galaxy overdensities. In this letter, we present another quadruple AGN candidate at including two SDSS quasars at a separation of roughly 480 kpc. The brighter quasar is accompanied by two AGN candidates (a type 1 AGN and a likely type 2 quasar) at close ( kpc) separation identified through emission line ratios, line widths and high ionization lines like N v . The extended Ly emission associated with the close triple system is more modest in extent and brightness compared to similar multiple AGN systems and could be caused by ram-pressure stripping of the type-2 quasar host during infall into the central dark matter halo. The predicted evolution of the system into a galaxy cluster with the AGN host galaxies forming the brightest cluster galaxy needs to be further tested by galaxy overdensity studies on large scales around the quadruple AGN candidate. If confirmed as a quadruple AGN with X-ray observations or rest-frame optical line ratios, this system would represent the second AGN quartet, the highest-redshift multiplet and the closest high-redshift triplet known.
@@
Galaxies: interactions – Galaxies: high-redshift – Galaxies: evolution – quasars: general – quasars: emission lines1 Introduction
@@
While single bright quasars may pinpoint protocluster cores, physically associated multiple active galactic nuclei (AGN) systems could be even more reliable tracers of local overdensities (Onoue et al., 2018; Sandrinelli et al., 2018), albeit also not a firm indicator (Fukugita et al., 2004; Green et al., 2011), as a number of highly active, massive galaxies require ample supply of cold gas feeding supermassive black holes (SMBHs) and star formation. In the framework of hierarchical structure formation, such objects are most likely to reside in peaks of the density field populating the rare massive end of the halo mass function (Bhowmick et al., 2020). In the absence of dense environments, multiple AGN systems would be exceedingly rare due to their transient nature (Hennawi et al., 2006). Interactions of three SMBHs are also predicted to be almost always necessary to form ultra-massive black holes with black hole masses (Hoffman et al., 2023), further highlighting their importance in cluster formation as the central brightest cluster galaxies commonly host ultra-massive black holes (e.g., Dullo et al. 2017; Mehrgan et al. 2019).
So far, only a few high-z multiple AGN systems on kiloparsec scales are known. These objects are likely in an early stage of galaxy merger and might share a dark matter halo, but not their host galaxies. In contrast to that, sub-arcsecond triple AGNs trace local late-stage mergers with multiple cores in the same host galaxy (e.g., Liu et al. 2011; Kalfountzou et al. 2017; Pfeifle et al. 2019). This work concentrates on the former AGN multiplets on kiloparsec scales, of which only five are known. Specifically, Hennawi et al. (2015) report the discovery of a quasar quartet at the peak of cosmic noon, . The AGN are embedded in a massive Ly nebula extending by more than 300 kpc, indicative of a substantial reservoir of cool ( K) gas (). The AGN are all found within a sphere of projected radius ¡150 kpc. The first triple quasar to be identified also resides at a redshift of 2 within a projected distance of less than 50 kpc, consistent with the typical onset of AGN activity in galaxy mergers (Djorgovski et al., 2007). Another triplet has been detected at , consisting of a close pair with a companion at 200 kpc projected distance (Farina et al., 2013). The third triplet known is likewise embedded in a massive Ly nebula at and comprises two quasars and one faint AGN at projected distances below 100 kpc (Arrigoni Battaia et al., 2018). A fourth triplet was found in a sample of almost 20,000 AGN candidates at a redshift of 1.13 (Assef et al., 2018).
Here, we report on the discovery of a quadruple AGN candidate at consisting of a close triplet, contained within 20 kpc in projection, and a forth quasar at a separation of 478 kpc. Throughout this work, we assume a flat CDM cosmology with , and (Planck Collaboration et al., 2020). Reported quasar magnitudes are in the AB system.
@@ 2 Observation and data analysis
One quasar in the system, SDSS J101254.73+033548.7 (hereinafter QSO1), has been targeted with the Multi Unit Spectroscopic Explorer (MUSE; Bacon et al. 2010) on the Very Large Telescope as part of a survey to study the environment of quasar pairs (Herwig et al., 2024). QSO1 is a known member of a quasar pair from the SDSS DR12 (Pâris et al., 2017). The projected distance to the second quasar in the pair SDSS J101251.06+033616.5 (hereinafter QSO2) is 478 kpc. For QSO2, only the SDSS spectrum is available. In SDSS DR17 (Abdurro’uf et al., 2022), the -band magnitude of the quasars integrated within the SDSS fiber with a diameter of 3 are mag and mag. Here we briefly describe the observations and data reduction. More detailed descriptions of sample selection, observation and subsequent data analysis can be found in Herwig et al. (2024). The MUSE data were acquired in Wide Field Mode on March 14, 2018 (seeing of ), yielding a field of view (FoV) of (pixel scale of ) and targeting only QSO1 as the separation of the pair is larger than the MUSE FoV. Three exposures were taken with on-source time each, rotated by 90 degrees with respect to each other and covering a wavelength range of with a sampling of . At the wavelength of Ly, the spectral resolution of corresponds to km s-1 ( km s-1 at C iii]). The data have been reduced following Farina et al. (2019) and González Lobos et al. (2023). Specifically, we use the MUSE pipeline v 2.8.7 (Weilbacher et al., 2014, 2016, 2020) to perform bias- and sky-subtraction, flat fields, twilight and illumination correction and wavelength and flux calibration. To remove residual sky emission lines, we use the Zurich Atmospheric Purge (Soto et al., 2016). After masking gaps between the CCDs and correcting offsets between the exposures with custom routines, the three exposures are median combined to obtain the final data cube. To take into account correlated noise introduced during data processing, the variance cube produced by the pipeline is rescaled layer-by-layer to the RMS of a background region in the science cube, resulting in a mean pixel-wise noise level at the position of QSO1 of .
@@ 3.1 Discovery of a candidate quadruple AGN
By collapsing the final MUSE cube along the wavelength axis, we obtained a continuum image of the field surrounding QSO1. In this image (r-band in Fig. 1, left), two additional continuum sources are visible in close proximity to QSO1. While similar geometries can be seen as a result of gravitational lensing of a single source, we can confidently reject this possibility in this case (Appendix A). We instead identify both of these sources (hereinafter called AGNc1 and AGNc2) as AGN candidates as detailed in the following. To constrain their nature, we extracted spectra from the MUSE cube within an aperture of 0.8 to minimize contamination from each other’s PSFs. To determine their positions, we first modeled the PSF of a star within the MUSE FoV with a 2D Moffat function and fit the model to the three continuum sources observed with MUSE (QSO1, AGNc1, AGNc2) by fixing the FWHM and leaving amplitude and position free to vary (Fig. 1, middle panel). The peak of the Moffat function was used as center of the aperture for spectral extraction and is indicated with pink circles (crosses) in the middle (right) panel of Fig. 1. To reduce noise while retaining the signal, the spectra were smoothed by a 1D Gaussian kernel with pixels, corresponding to . Before smoothing the spectrum of AGNc2, we masked a sky emission line at 6365.3 as it would otherwise be convolved with the C iv line (Appendix E). The spectra are shown in Fig. 2, together with the SDSS spectrum of QSO2. The two QSOs and two AGN candidates are all detected in Ly, confirming that they reside at roughly the same redshift. AGNc1 tentatively shows a broad C iii] line and has excess flux in N v , Si iv and He ii , while in the spectrum of AGNc2, there is a tentative detection of a narrow C iii] and He ii line as well as a prominent narrow, blue-shifted C iv line. Within the aperture of used for spectral extraction, the -band magnitude for the two AGN candidates is mag and mag111In this smaller aperture, the magnitude of QSO1 decreases to mag. Between QSO1 and AGNc1, we measured an angular separation of 2.1, corresponding to the distance kpc at this redshift. QSO1 and AGNc2 are separated by 2.6 or 20 kpc, and QSO2 is 61.6 or 478 kpc away from QSO1. We obtained the systemic redshift by fitting the C iii] line with a double Gaussian model and using the peak of the line complex as reference wavelength. The value obtained for AGNc1 is likely overestimated as a redshift of implies a highly unusual Ly line shape dominated by the blue wing (Fig. 4). If this was the case, the He ii line might be misidentified and possibly not detected. Positions of the AGN candidates and projected distance between each other as well as redshift and velocity shifts are summarized in Table 1.
@@
We fit the spectra as described in Appendix B and summarize the resulting line fluxes, widths and upper flux limits in Table 2. The line width of AGNc2 is narrow with km s-1 and consistent with a Type II AGN, while the spectrum of AGNc1 is dominated by broad lines of km s-1 indicative of a Type I AGN.
3.2 Properties of QSO1 and QSO2
@@
Using the fitting results (Appendix B), we can derive additional properties for the brighter SDSS quasars QSO1 and QSO2. We calculated the black hole masses using the relation and absolute accuracy estimate derived for C iv in Vestergaard & Peterson (2006) and obtained and . At the position of QSO1, radio emission is detected with a flux density of in VLASS (Lacy et al., 2020) and in VLA FIRST (Becker et al., 1994), leading to a radio slope of (). Additionally using the UV continuum slope from the fit to the spectrum of QSO1 (Appendix B) of (), we calculated the ratio (Kellermann et al., 1989) by extrapolating the rest-frame flux densities and obtained , classifying QSO1 as radio-loud. Although no additional radio emission is detected, we cannot exclude that AGNc1 and AGNc2 are radio-loud due to the shallow surveys depth222The VLASS continuum image rms is Jy/beam (Lacy et al., 2020), FIRST data has an rms of Jy/beam close to the system location(Becker et al., 1994). and their dim UV continuum.
@@
3.3 Emission line diagnostics
A commonly adopted diagnostic diagram to distinguish AGNs from star forming galaxies (SFGs) in rest-frame UV uses the line ratios C iv /C iii] (C43), mainly influenced by the ionization parameter, and C iii]/He ii (C3He2), which is sensitive to metallicity and allows distinction between shocks and AGN photoionization (Feltre et al., 2016). Hirschmann et al. (2019) propose demarcation lines to distinguish between active and inactive galaxies, but current models are not capable of fully reproducing observed line ratios, leading to large impurities in the selection of AGNs and specifically faint AGNs. Classification is further complicated because certain types of galaxies would not appear in such a diagnostic diagram: high- radio galaxies (HzRGs) commonly have no detection of C iv (56 % in De Breuck et al. 2000) and SFGs usually show C iv and Si iv in absorption instead of emission (e.g. Shapley et al. 2003; Calabrò et al. 2022). In Fig. 3 we show the line ratios of the candidate AGNs and SDSS quasars together with line ratios of SFGs and different types of AGNs from the literature. Both quasars and both AGN candidates as well as 30 % of the literature AGNs fall within the composite region of the diagram, indicating contribution from both star formation and AGN activity. Surprisingly, the line ratios of QSO1 indicate the highest contribution from star formation of all four objects although it is definitively classified as quasar, further highlighting the difficulty in AGN selection based on UV line ratios. It is also notable that the stacks of observed SFGs selected based on C iii] emission are classified as composite objects as well.
However, the detection of N v and Si iv , two highly ionized lines, and the extreme line width of AGNc1 larger than firmly point toward an AGN origin. If these broad lines were produced by galactic-scale outflows instead, the violent kinematics should lead to a high contribution from shocks and likely lead to a detectable C iv emission line. The upper limit on the EW of C iv lies above the AGN demarcation line of at the corresponding C iv /He ii ratio limit for AGNc1 of ¡ 1 (Nakajima et al., 2018) and AGN ionization can therefore not be excluded based on the non-detection of the line. The ratio of Si iv to He ii exceeds those typical for HzRGs (Humphrey et al., 2008) as well as quasars (Nagao et al., 2006). On the other hand, the C iv EW of AGNc2 lies firmly above the demarcation line of 12 Å in Nakajima et al. (2018) at the corresponding C iv /He ii ratio for AGNc2, indicating that this source is an AGN.
@@ 4.1 Clues from the peculiar extended Ly emission
The pink contours in Fig. 1 show the 2 and 5 levels of extended Ly emission, corresponding to a surface brightness of and respectively, and determined after subtracting empirical quasar point spread functions and continuum sources from the combined cube. This analysis is described in detail in Herwig et al. (2024) and González Lobos et al. (2023). The nebular emission extends for about 50 kpc in the south-east direction with respect to QSO1, starting from the projected location of AGNc2. Additional Ly emission coincides with the position of AGNc1. Surprisingly, the surface brightness level of the extended Ly emission is much lower than typically found for multiple AGN systems (Hennawi et al., 2015; Arrigoni Battaia et al., 2018) and the AGN candidates are not embedded in the nebula. Instead, the flux-weighted centroid of the extended emission is offset from QSO1 by 34 kpc (Herwig et al., 2024). Possible explanations for this could be a faster evolution and therefore early cool gas depletion of outlier density peaks in the universe or a later merger stage of the close AGN triplet candidate presented in this work. Moreover, strong feedback effects might have already depleted the circumgalactic medium of this system of cool gas.